OpenStudy (chris215):

Find the area of the region bounded by the graphs of y = x, y = -x + 4, and y = 0.

OpenStudy (sshayer):

 or\[reqd.~area=\int\limits_{0}^{2}x dx+\int\limits_{2}^{4}\left( -x+4 \right)dx=?\]